OpenStudy (anonymous):

dy=(x^4-1)/(3x) Find deriavative??

OpenStudy (primeralph):

dy or y?

OpenStudy (dumbcow):

use quotient rule \[(\frac{f}{g})' =\frac{f'g -fg'}{g^{2}}\]

OpenStudy (zzr0ck3r):

\[y=\frac{1}{3}x^3+\frac{1}{3}x^{-1}\]

OpenStudy (anonymous):

sorry its not dy its y

OpenStudy (zzr0ck3r):

\[\frac{d(ax^n)}{dx}=nax^{n-1}\]is the only thing you need to use

OpenStudy (anonymous):

use it

OpenStudy (zzr0ck3r):

what is the derivative of \[\frac{1}{3}x^3\]?

OpenStudy (zzr0ck3r):

@fakharabbas786 we are not here to do it for you, we are here in hopes that we can show you how to do it.

OpenStudy (zzr0ck3r):

do you know the derivative of x^3?

OpenStudy (anonymous):

3x^2

OpenStudy (zzr0ck3r):

so what is the derivative of (1/3)x^3

OpenStudy (zzr0ck3r):

its the same thing multiplied by (1/3)

OpenStudy (zzr0ck3r):

what is it then?

OpenStudy (anonymous):

its confusing me

OpenStudy (zzr0ck3r):

so when we take the derivative of something like (1/3)x^3 we take the derivative of x^3 and as you said this is 3x^2 but we still have the (1/3) out front. so its (1/3)*3*x^2 = x^2

OpenStudy (zzr0ck3r):

\[3\frac{1}{3}x^2\]

OpenStudy (zzr0ck3r):

so what is the derivative of \[\frac{1}{3}x^{-1}\]?

OpenStudy (zzr0ck3r):

same thing as before, we bring down the -1 and then de-increment the exponent by 1 \[(-1)\frac{1}{3}x^{-2}=\frac{-1}{3}x^{-2} = \frac{-1}{3x^2}\]

OpenStudy (anonymous):

zzrock3r u r confusing me

OpenStudy (zzr0ck3r):

\[so\space the\space derivative\space of\\\frac{x^4-1}{3x}=\frac{1}{3}x^3+\frac{1}{3}x^{-1}\\is\\x^2-\frac{1}{3x^2}\]

OpenStudy (zzr0ck3r):

first we take the derivative of (1/3)x^3 and get x^2 then we take the derivative of (1/3)x^(-1) and get (1/(3x^2))

OpenStudy (zzr0ck3r):

then we add them together

OpenStudy (zzr0ck3r):

tell me what part is confusing

OpenStudy (zzr0ck3r):

you two are lame.....